the average temperature last may was 80 degrees and the average temperature last june was 90 degree. what percent did the average temperature increase from may to june?
1
Expert's answer
2019-09-18T11:48:53-0400

"i=\frac{90-80}{80}=\frac{1}{8}=" 0.125=12.5%